You are considering enrolling in a 2-year long on-line MBA program. You will continue your current employment. If you do, you must pay two payments of $30,000; the first happens today, at the beginning of the program, and the second is due 1 year from today, at the beginning of the second year. These payments are tax deductible since the program qualifies as education necessary for your job. Because of this advanced degree, you expect your annual earnings stream will increase; this increase will occur after you complete the 2-year program. You plan on working for 20 years after you graduate. Remember that in capital budgeting, we assume that cash flows from the operations/activity happen at the end of the year. To finance this program, you will withdraw money from your investment portfolio, which is currently earning an average of 6%. Thus, you have decided that any use of these funds should earn at least 6%. What is the annual, before-tax increase in earnings necessary to make this investment worthwhile? Your personal tax rate is 25%.
Answer #1

To make this investment worthwhile, the present value of after-tax cash inflows should be higher than the present value of after-tax cash outflows.

Present value = future value / (1 + interest rate)number of years

The first cash flow is at year 0, and the second is at year 1.

After-tax cash outflow = cash outflow * (1 - tax rate)

After-tax cash outflow = $30,000 * (1 - 25%) = $22,500

Present value of after-tax cash outflows = ($22,500 / (1 + 6%))0 + ($22,500 / (1 + 6%))1

Present value of after-tax cash outflows = $43,726.42.

The present value of after-tax increase in earnings should be at least $43,726.42.

The after-tax increase in earnings is calculated using PMT function in Excel :

rate = 6%

nper = 20 (20 years of increased earnings)

pv = -43726.42 (required present value. This is entered with a negative sign because it is like the amount to purchase an annuity)

PMT is calculated to be $3,812.27

A4 6 A contok 20. 70 foc =PMT(6%,20,-43726.42) 4 $3,812.27

Before-tax cash inflow = after-tax cash inflow / (1 - tax rate)

Before-tax cash inflow = $3,812.27 / (1 - 25%)

Before-tax cash inflow = $2,859.20.

The annual, before-tax increase in earnings necessary to make this investment worthwhile is $2,859.20.